### v1

# Small Stickers, Big Meanings: A Multilingual Sticker Semantic Understanding Dataset with a Gamified Approach

Abstract.
Stickers, though small, are a highly condensed form of visual expression, ubiquitous across messaging platforms and embraced by diverse cultures, genders, and age groups. Despite their popularity, sticker retrieval remains an underexplored task due to the significant human effort and subjectivity involved in constructing high-quality sticker query datasets. Although large language models (LLMs) excel at general NLP tasks, they falter when confronted with the nuanced, intangible, and highly specific nature of sticker query generation.
To address this challenge, we propose a threefold solution. First, we introduce Sticktionary, a gamified annotation framework designed to gather diverse, high-quality, and contextually resonant sticker queries. Second, we present StickerQueries, a multilingual sticker query dataset containing 1,115 English and 615 Chinese queries, annotated by over 60 contributors across 60+ hours. Lastly, Through extensive quantitative and qualitative evaluation, we demonstrate that our approach significantly enhances query generation quality, retrieval accuracy, and semantic understanding in the sticker domain. To support future research, we publicly release our multilingual dataset [^1] https://huggingface.co/datasets/metchee/sticker-queries along with two fine-tuned query generation models [^2] https://huggingface.co/metchee/sticker-query-generator-en [^3] https://huggingface.co/metchee/sticker-query-generator-zh.
Resource, Sticker Retrieval, Sticker Query Generation.
journalyear: 2025conference: Under Review.; ;
1. Introduction
Stickers have become an integral component of modern digital communication. Widely used across diverse demographics—spanning genders, cultures, and age groups—they enrich messaging experiences by conveying emotions, reactions, and humor in a compact, expressive format. While several sticker datasets have been developed, few resources focus on the task of sticker retrieval, despite its centrality in real-world applications. This gap is largely due to the intensive human effort required to create high-quality datasets, and the challenges of interpreting and formalizing the rich semantics embedded in sticker use.
Despite their small size, stickers are capable of conveying deeply nuanced human emotions. For instance, expressing mild annoyance, frustration, or sarcasm might be easily achieved with a single sticker (e.g., Fig. 1), but would require a verbose textual explanation. Sticker queries often involve a rich interplay of modalities—such as sound (e.g., “bleh bleh bleh”), action (“ghost face”), emotion (“whatever”), culture (“nenenipoopoo”), thought (“can’t do nothing”), and tone. These phenomena are observed across both Chinese and English contexts. Generating such queries requires cultural sensitivity, contextual understanding, and linguistic precision—capabilities that even advanced vision-language models (VLMs) struggle to replicate. Sticker query generation is prone to hallucinations, and typical models often fail to grasp the subtle social and emotional layers stickers encode.
Figure 1. Base VLM models are unable to understand and generate accurate sticker search queries, which requires depth and precision, oftentimes hallucinating. This persists in both English and Chinese languages.
To address these limitations, we introduce StickerQueries, a multilingual sticker query dataset for understanding and retrieving stickers in context. We summarize our contributions in the following:
  1. (1)
    We propose Sticktionary, a gamified annotation framework that encourages contributors to create expressive, context-rich sticker queries through an engaging game-based interface. This design fosters the generation of high-quality and semantically resonant queries.
  2. (2)
    We construct and release StickerQueries, a dataset consisting of 1,115 English and 615 Chinese queries, annotated by over 60 contributors with more than 60 hours of effort. We also release two fine-tuned query generation models on Hugging Face.
  3. (3)
    Through extensive experiments, we demonstrate significant improvements in sticker retrieval and query generation performance, underscoring the effectiveness of our dataset and methodology.
2. Related Work
Table 1. Comparison of sticker query datasets sorted by year. Datasets supporting English queries and publicly available are highlighted in bold. Until now, no sticker intention dataset has been publicly available except our proposed StickerQueries.
Dataset Year Description Pub. Avail? Have Queries? Query Language? Native English Query?
Sticker Dialogues
SRS, PESRS (Gao et al., 2021, 2020) 2020 Sticker dialogues.
MOD (Fei et al., 2021) 2021 Sticker dialogues
StickerInt (Liang et al., 2024) 2024 Sticker dialogues
MCDSCS (Shi and Kong, 2024) 2024 Sticker dialogues.
STICKERCONV(Association for Computational Linguistics 2024 and Zhang, 2024) 2024 Simulated sticker dialogues.
U-Sticker (Chee et al., 2025) 2025 Continuous sticker dialogues.
MultiChat (Wang et al., 2025) 2025 Sticker dialogues.
Sticker Sentiment Analysis
CSMSA (Ge et al., 2022) 2022 Sticker sentiment.
Sticker820K (Zhao et al., 2023) 2023 Sticker sentiment.
SER30K (Liu et al., 2022) 2022 Sticker sentiment.
Basic Sticker Collections
ChineseB2B (Zhaoolee, 2024) 2000 Sticker images.
TGIF (Li et al., 2016) 2016 GIF captioning.
VSD2M (Yuan et al., 2025) 2025 Animated sticker dataset.
Sticker Queries
StickerTag (Wang et al., 2024) 2024 Sticker-tag. Unknown Unknown Unknown
PerSRV (Chee et al., 2024) 2024 Sticker search query. ZH
StickerQueries 2025 Sticker search query. EN, ZH
With the onset of instant messaging, we have observed significant advancements in sticker dataset retrieval. However, given the complexity and diversity of sticker tasks, existing datasets still represent only the tip of the iceberg. We broadly categorize these datasets into four groups: sticker dialogues, sticker sentiment analysis, basic sticker collections, and sticker queries.
Sticker dialogues typically consist of truncated conversations designed to support sticker retrieval or recommendation (Gao et al., 2021, 2020; Fei et al., 2021; Liang et al., 2024; Shi and Kong, 2024; Association for Computational Linguistics 2024 and Zhang, 2024; Wang et al., 2025), whereas U-Sticker (Chee et al., 2025) uniquely provides continuous user information, temporal context, and domain-specific sticker conversations. Among these, several datasets are publicly available, while others remain private or unavailable.
Sticker sentiment datasets focus on predicting the sentiment conveyed by stickers, either as binary positive/negative labels (Ge et al., 2022) or more fine-grained classifications (Liu et al., 2022). Some, like Sticker820K (Zhao et al., 2023), also include supplementary information such as OCR text and descriptive metadata. Within this category, SER30K is publicly accessible, while others are not.
Basic sticker collections mainly provide raw sticker images or GIFs for general usage or generation tasks. Examples include ChineseB2B (Zhaoolee, 2024), which offers a large collection of sticker images, TGIF (Li et al., 2016) for GIF captioning, and the recently introduced VSD2M (Yuan et al., 2025), an animated sticker dataset intended for sticker generation.
Finally, sticker queries datasets are critical as they capture the underlying intention behind sticker usage through condensed, highly precise keywords or search queries. Unfortunately, datasets like StickerTag (Wang et al., 2024) and PerSRV (Chee et al., 2024)are not publicly available for analysis. Though PerSRV (Chee et al., 2024) provides an open source model, their training is limited purely to Chinese queries, restricting its applicability to English-language research. With contrasting cultural nuances, naive translation between languages fails to adequately capture sticker usage intentions. Our StickerQueries dataset addresses these gaps by providing bilingual (English and Chinese) queries crafted by native speakers, thereby bridging the semantic understanding of stickers across diverse cultural contexts and enabling more effective retrieval and recommendation systems.
2.1. Sticker Semantic Understanding
2.2. Image Captioning
Recent models like BLIP and BLIP-2 (Li et al., 2022, 2023a), as well as GIT (Wang et al., 2022a) and OFASys (Wang et al., 2022b), have advanced image captioning through large-scale vision-language pretraining and unified generative modeling. VinVL (Zhang et al., 2021) and PaLI (Chen et al., 2023) further improve visual representations and multilingual capabilities. While these models produce accurate and fluent descriptions, they focus mainly on tangible visual semantics. In contrast, sticker search revolves around emotional expression, which traditional captioning models are not designed to capture, making them inadequate for emotion-driven query generation.
2.3. Large Language Model for Sticker Semantic Understanding
Large Language Models (LLMs), such as ChatGPT, LLaMA, and their multimodal extensions (Achiam et al., 2024; Chang et al., 2024; Touvron et al., 2023; Alayrac et al., 2022; Li et al., 2022, 2023b; Zhu et al., 2023; Liu et al., 2023; Dai et al., 2024; Wang et al., 2023), have shown strong capabilities in integrating visual and textual information. Models like Flamingo and InstructBLIP introduce cross-modal alignment mechanisms, while BLIP, MiniGPT, and LLaVA variants connect frozen vision encoders with LLMs to enable richer multimodal understanding. These models hold promise for sticker semantics, where emotional and contextual cues are key. However, vanilla LLMs still struggle to generate intent-aligned sticker queries. PerSRV (Chee et al., 2024) demonstrates this gap, showing that supervised fine-tuning is needed but its robustness across languages remains largely questionable. This highlights the necessity for a sticker query dataset which can improve sticker understanding and retrieval tasks.
2.4. Games with a Purpose (GWAP)
GWAP leverages human engagement through gameplay (von Ahn and Dabbish, 2008) and the low-pressure environment makes it particularly effective in gathering emotionally rich or subjective data which is often difficult to obtain through traditional annotation pipelines. Methodologies include output-agreement games, games, where players independently try to generate the same label for an input, encouraging consensus on common interpretations; inversion-problem games,, where one player describes an item and the other guesses it, capturing descriptive and emotionally resonant cues; and input-agreement games, where players judge whether they received the same input, supporting perceptual or semantic similarity annotations (Law and von Ahn, 2009). GWAPs are especially effective and enjoyable, especially when combined with engaging visual effects, to enhance data quality. This makes them ideal for the sticker annotation task.
3. Motivation and Goals
3.1. Sticker Usage Behavior Survey
To better understand user sticker usage behavior, we conducted a survey covering variables such as age, primary language, average number of text messages sent before a sticker is used, the number of stickers used per day, and the motivations behind sticker usage.
We collected a total of 192 responses from a linguistically diverse user base and present our findings here. Interestingly, both genders reported the same average daily sticker usage—approximately 7.55 stickers per day. While younger users generally tend to send more stickers, users across all age groups expressed genuine enjoyment in using them. On average, participants reported sending a sticker every four text messages. Notably, none of the 192 respondents reported zero sticker usage, underscoring the ubiquity of stickers in digital communication.
Figure 2. Emotional expression emerged as the most common reason for using stickers, while other motivations—such as actions, greetings and thoughts—were mentioned with relatively equal frequency.
As shown in Fig 2, the most common reason for using stickers is emotional expression, such as conveying feelings like happiness, sadness, or confusion. While this was the most dominant purpose cited by participants, other motivations were mentioned with relatively equal frequency. Users also mentioned using stickers for politeness or to clarify tone. These diverse uses reflect the rich communicative potential of stickers, not only do stickers enhance conversations but also make digital interactions more expressive, nuanced, and enjoyable. In light of these insights, it becomes clear that stickers play a pivotal role in digital communication—used frequently, across age groups, and for a wide range of expressive functions. The universality and versatility of sticker usage reveal a communication tool that is both culturally rich and deeply embedded in everyday interaction. With a vast and engaged user base, and an array of meaningful use cases, further exploration into sticker behavior holds tremendous potential—not only to advance human-computer interaction but to enrich how we understand and design for user expression.
3.2. Dataset Goals
With this motivation in mind, we intend to develop the first multilingual sticker queries dataset publicably available. We establish the following goals.
  • FAIR Compliance (Wilkinson et al., 2016): Ensure the dataset is openly accessible, well-documented, and licensed for reproducibility and reuse.
  • Natural Associations: Focus on commonly used stickers and intuitive textual expressions, avoiding overly niche or esoteric cases.
  • Multilingual Coverage: Move beyond direct translations to include culturally and contextually appropriate expressions.
  • High Quality Representative Queries: Capture expressions used by native speakers, including local slang and region-specific terms (e.g., “savage”).
  • Diverse Expressions: Reflect varied phrasing for similar intents (e.g., “haha,” “lol,” “lmao”).
  • Resonant Queries: Ensure queries are broadly understandable and relatable, not limited to a narrow subgroup.
4. Sticktionary: Game-Based Sticker Semantics Annotation Framework
4.1. Game Design
Figure 3. Sticktionary: A game-based framework for sticker semantics annotation. Players are randomly assigned initial roles and alternate between labeling and retrieving—annotating stickers with textual hints and identifying stickers based on those hints. Tasks are drawn from a curated pool and may be skipped. Annotated stickers enter a temporary retrieval pool, and successful retrievals mark task completion.
Notably, collecting high-quality representative queries, diverse expressions, and broadly resonant phrases is particularly challenging in traditional annotation pipelines. Agreeableness is especially difficult to model since sticker search is subjective, and ensuring high-quality representation often demands significant annotator effort and attention, potentially prolonging the annotation process. To address these challenges, we leverage the engaging and expressive visual modality of stickers and propose a game-based annotation framework, Sticktionary, illustrated in Fig 3, is inspired by the GWAP paradigm. Sticktionary features two alternating roles—labeler and retriever—designed to collect and evaluate sticker search queries through interactive and enjoyable gameplay.
Labeler In each round, the labeler is shown a sticker and asked imagine themselves searching for the sticker, in which they they are expected to key in search queries. Then, their queries are used to search and a preliminary sticker result is shown. This imaginative state prevents player from merely describing the features of the sticker. Based on this, they may revise their queries accordingly. Players are also instructed to avoid repetitive search queries, and purely visual descriptions.
Retriever The retriever then receives the labeler’s queries and must identify the correct sticker from a set of candidates, selecting and ranking up to three. Additionally, retrievers may optionally contribute new query suggestions for any sticker they feel confident about, encouraging the organic collection of emotionally resonant or imaginative queries. The outcome of the retrieval (successful or not) is shared with the labeler as feedback to refine future queries. Should the players feel that either task is too challenging they can choose to skip.
Ensuring Quality Sticktionary’s alternating roles naturally sustain attention and reduce fatigue, promoting higher quality annotations. The point-based system turns precision into reward which further motivates players to craft high quality annotations.
Ensuring Resonance Intuitively, if one player’s query guides another to the correct sticker, it’s a strong indicator of shared understanding and a resonant query. In addition, real-time feedback loops nudge players toward phrasing that feels universally intuitive, rather than overly niche or idiosyncratic.
Ensuring Diversity Lastly, the involvement of more than one round of annotation can ensure diversity and variety. Likewise, if queries can be effectively retrieved, their quality can be ensured.
These mechanisms function in tandem rather than isolation. The built-in game design’s role-switching, feedback loops, and incentive structures collectively ensure that the collected annotations are not only of high quality, but also diverse and widely resonant.
4.2. Game Settings
4.2.1. Player Recruitment
We recruit participants by incentivizing top performers with themed prizes featuring the university’s mascot panda, along with a participation reward of 30 RMB for completing an observable annotation session.
4.2.2. Data Selection
To ensure the goal of natural sticker associations, we select messages preceded by at least twenty words of dialogue, filtering out command-driven interactions and exchanges dominated by brief replies. This ensures that the stickers appear in expressive, emotionally grounded contexts. We apply these filters to the U-Sticker dataset (Chee et al., 2025), followed by a manual review. The final collection consists of 1,136 non-duplicated conversations, each curated for contextual richness and intuitive alignment.
4.3. Game Statistics
Table 2. Summary of player participation and task distribution across English and Chinese versions of the game after outlier removal. The dataset reflects contributions from over 62 players and approximately 60 hours of gameplay.
Metric English Chinese
# Label-role players 42 18
# Retrieval-role players 43 17
Avg. tasks per label user 61.1 78.7
Avg. tasks per retrieve user 55.0 68.5
In total, we observed 60 hours worth of gameplay sessions which involved over 40 English players and 22 Chinese players. Table 3 provides a comprehensive snapshot of participant engagement and workload distribution across labeling and retrieval activities.
For successfully retrieved samples, we append them directly. Failures undergo manual review, where most ill retrievals stem from a flood of semantically similar stickers—making retrieval results actually reasonable and acceptable. Inaccurate queries are discarded. Lastly, obvious spelling errors are corrected.
5. StickerQuery
5.1. Statistics
Table 3. StickerQuery comprises over 1,700 multilingual sticker-query pairs, each annotated by at least two annotators, offering a rich, diverse set of natural expressions.
Metric English Chinese
Unique sticker query pair 1,116 615
Unique terms 5,347 1,944
Avg. # queries per sticker 6.77 4.60
Table 3 presents overview of the StickerQuery dataset statistics, highlighting its scale, linguistic diversity, and annotation coverage. For the following steps, we use Jieba for Chinese tokenization (Junyi, 2012).
Figure 4. Word cloud of the English StickerQueries, feelings dominate the visual.
Figure 5. Frequency distribution of the top 10 most common English query words in the StickerQueries dataset.
We find that emotional expressions dominate sticker queries: 9 of the top 10 entries in both English and Chinese pertain to feelings rather than actions or visuals (Figures 5 and 7). This aligns with the idea that users are often driven by internal emotional states when searching, rather than external depictions.
Interestingly, although survey participants emphasized action-oriented queries, annotation patterns reveal a stronger inclination toward emotional resonance—affirming the sticker’s role as an affective medium. While English and Chinese users show largely parallel emotional vocabularies, terms like lol, chill, and zen emerge as uniquely English, suggesting subtle cultural inflections in how emotion is expressed across languages.
Figure 6. Word cloud of the Chinese StickerQueries, feelings dominate the visual.
Figure 7. Top 10 most common Chinese query words in the StickerQueries dataset, showing the expressions trends.
5.2. Interannotator Similarity
Table 4. Multilingual sticker query similarity evaluation. BLEU and ROUGE are low due to short queries, while BERT-based and cosine similarity show higher semantic alignment.
Metric En Zh
BLEU (Papineni et al., 2002): 0.0756 0.0132
ROUGE-2: 0.1524 0.0034
ROUGE-L: 0.2796 0.0195
CosineSim: 0.7131 0.7728
BERTScore (P) (Zhang et al., 2020) 0.6710 0.7061
BERTScore (R) 0.6666 0.6983
BERTScore (F1) 0.6628 0.6986
We use the pretrained BERT-base models released by Google for English (Devlin et al., 2019) and Chinese to compute both cosine similarity and BERTScore (Zhang et al., 2020) between query annotations. ROUGE-1, ROUGE-2, and ROUGE-L (Lin, 2004) are also reported for comparison with traditional string-based metrics.
5.2.1. English Annotations
English annotations show higher lexical overlap, with BLEU (0.0756) and ROUGE scores (e.g., ROUGE-1: 0.3019), indicating strong n-gram similarity. Semantic metrics also confirm good meaning alignment (cosine: 0.7131, BERT F1: 0.6628).
5.2.2. Chinese Annotations
Chinese annotations score lower on lexical metrics (BLEU: 0.0132, ROUGE-1: 0.0195), likely due to greater paraphrasing. However, they achieve higher semantic similarity (cosine: 0.7728, BERT F1: 0.6986), indicating strong meaning preservation despite fewer surface overlaps.
Following prior work (Zhang et al., 2020), a BERTScore F1 above 0.65 indicates moderate semantic agreement in open-ended generation tasks. Our dataset achieves 0.66–0.70, validating the reliability of annotations while retaining expressive diversity.
6. Applications
6.1. Query Generation
We split the data into 20% test and 80% training sets and fine-tuned LLaVA-1.5B-7B, a moderately sized vision-language model, on our StickerQueries dataset.
Table 5. Performance of LLaVA-1.5-7B on StickerQueries, showing zero-shot vs fine-tuned results. Significant improvements demonstrate effective adaptation to both English and Chinese query generation.
Model BLEU ROUGE-2 ROUGE-L CosSim
English Evaluation
Zero-Shot 11.08 2.82 19.22 0.778
Fine-Tuned (SFT) 15.62 8.26 27.05 0.811
Improvement $\uparrow$40.9% 192.5% 40.8% 4.2%
Chinese Evaluation
Zero-Shot 0.5138 0.0296 2.2584 0.6349
Fine-Tuned (SFT) 6.3432 2.970 24.26 0.7976
Improvement $\uparrow$1134.6% 9934.5% 974.2% 25.6%
Fine-tuning our StickerLabel dataset leads to clear improvements across all evaluation metrics for both English and Chinese, in particular significant gains in ROUGE scores and cosine similarity. This demonstrates the dataset’s effectiveness in enhancing the model’s ability to generate accurate, contextually relevant sticker queries.
6.2. Sticker Retrieval
Table 6. Recall@K performance comparison across query generation methods.
Model English Chinese
R@1/50 $\Delta\%$ R@1/50 $\Delta\%$
Context 0.018/0.297 900/129 0.008/0.407 413/12
BLIP2 0.032/0.248 462/174 0.005/0.225 720/102
Vanilla VLM 0.153/0.644 18/6 0.016/0.423 156/8
SFT VLM 0.180/0.680** 0.041/0.455**
Since we used a 20% test split, we compare three sticker semantic baseline methods for the sticker retrieval task to generate queries.
  • Context: Take the previous utterance in the context as the query.
  • BLIP2 (Li et al., 2023a): Generate queries via image captioning.
  • Vanilla VLM(Liu et al., 2023): Zero-shot query generation with LLaVA-1.5B-7B.
  • SFT VLM (Ours): Finetuned LLaVA-1.5B-7B model on StickerQueries dataset (Liu et al., 2023).
We generate test candidate pools using the above baselines and use BM25 (Robertson et al., 2009) for retrieval. The fine-tuned LLaVA model trained on the StickerQueries dataset achieves the highest Recall@K scores at every cutoff, indicating it generates more accurate and relevant queries for sticker retrieval. Compared to zero-shot baselines like Vanilla VLM, BLIP2 image captions, and direct conversation context, our model provides a significantly stronger candidate pool for BM25 retrieval, underscoring the benefits of domain-specific fine-tuning.
7. Discussion and Future Work
We understand that our dataset is limited and expanding it would likely enhance performance and robustness. Additionally, incorporating more multilingual data and increasing the number of survey respondents could help us to understand sticker usage behavior.
8. Conclusion
In this work, we propose StickerQueries, a bilingual benchmark designed to support and evaluate sticker understanding and retrieval in realistic conversational scenarios. By leveraging a gamified annotation process, we build a dataset that captures rich, high-quality queries reflecting the subtle and often culturally grounded semantics of sticker use. Our experiments show that fine-tuning vision-language models on this dataset leads to substantial performance improvements in both query generation and retrieval tasks. These findings highlight the limitations of general-purpose models and the need for specialized datasets in high-context, multimodal applications. We believe StickerQueries provides a valuable foundation for future work in personalized visual communication, semantic understanding, and cross-lingual sticker interaction.